# Large Language Monkeys00footnotetext: Title inspired by https://en.m.wikipedia.org/wiki/Infinite_monkey_theorem.: Scaling Inference Compute with Repeated Sampling

### v2

@@ Abstract
Scaling the amount of compute used to train language models has dramatically improved their capabilities. However, when it comes to inference, we often limit the amount of compute to only one attempt per problem. Here, we explore inference compute as another axis for scaling by increasing the number of generated samples. Across multiple tasks and models, we observe that coverage – the fraction of problems solved by any attempt – scales with the number of samples over four orders of magnitude. In domains like coding and formal proofs, where all answers can be automatically verified, these increases in coverage directly translate into improved performance. When we apply repeated sampling to SWE-bench Lite, the fraction of issues solved with DeepSeek-Coder-V2-Instruct increases from 15.9% with one sample to 56% with 250 samples, outperforming the single-attempt state-of-the-art of 43% which uses more capable frontier models. Moreover, using current API pricing, amplifying the cheaper DeepSeek model with five samples is more cost-effective and solves more issues than paying a premium for one sample from GPT-4o or Claude 3.5 Sonnet. Interestingly, the relationship between coverage and the number of samples is often log-linear and can be modelled with an exponentiated power law, suggesting the existence of inference-time scaling laws. Finally, we find that identifying correct samples out of many generations remains an important direction for future research in domains without automatic verifiers. When solving math word problems from GSM8K and MATH, coverage with Llama-3 models grows to over 95% with 10,000 samples. However, common methods to pick correct solutions from a sample collection, such as majority voting or reward models, plateau beyond several hundred samples and fail to fully scale with the sample budget.
00footnotetext: Equal Contribution. Work done by BB as a visiting researcher at Stanford.
@@ 1 Introduction
The ability of large language models (LLMs) to solve coding, mathematics, and other reasoning tasks has improved dramatically over the past several years [46, 11]. Scaling up model training has been a consistent driver of these gains. Investments in larger models, larger pre-training datasets, and more extensive post-training (e.g. through collecting human preference labels) has led to remarkably capable generalist systems [2, 3, 4, 51].
In contrast, a comparatively limited investment has been made in scaling the amount of computation used during inference. Larger models do require more inference compute than smaller ones, and prompting techniques like chain-of-thought [61] can increase answer quality at the cost of longer (and therefore more computationally expensive) outputs. However, when interacting with LLMs, users and developers often restrict models to making only one attempt when solving a problem. In this work, we investigate repeated sampling (depicted in Figure 1) as an alternative axis for scaling inference compute to improve LLM reasoning performance.
Figure 1: The repeated sampling procedure that we follow in this paper. 1) We generate many candidate solutions for a given problem by sampling from an LLM with a positive temperature. 2) We use a domain-specific verifier (ex. unit tests for code) to select a final answer from the generated samples.
@@
    Precision: In settings where we must select a final answer from the collection of generated samples, can we identify correct samples?
With unlimited attempts, any model that assigns a non-zero probability to every sequence will achieve perfect coverage. However, repeated sampling is only practical if we can improve coverage with a feasible budget. Moreover, without the ability to decide between samples, the applications of repeated sampling are limited. Existing work provides encouraging evidence along both of these directions, showing examples of repeated sampling improving performance in math, coding, and puzzle-solving settings [60, 47, 23]. Notably, AlphaCode [40], a state-of-the-art system for competitive programming, finds that performance continues to improve with a million samples per problem.
@@
Here, we show that repeated sampling is an effective methodology for improving coverage across a range of tasks, models, and sample budgets. For example, when solving CodeContests [40] programming problems using Gemma-2B [52], we increase coverage by over 300x as we scale the number of samples, from 0.02% with one attempt to 7.1% with 10,000 attempts. Interestingly, the relationship between $\log(\text{coverage})$ and the number of samples often follows an approximate power law. With Llama-3 [3] and Gemma models, we observe that coverage grows nearly log-linearly with the number of samples over several orders of magnitude.
@@
In settings where all of a model’s solutions can be automatically verified, such as with proof checkers or unit tests, these increases in coverage translate directly into improved task performance. When applying repeated sampling to competitive programming and writing Lean proofs, models like Llama-3-8B-Instruct can exceed the single-attempt performance of much stronger ones like GPT-4o [2]. This ability to amplify weaker models extends to the challenging SWE-bench Lite dataset of real-life GitHub issues [32], where the current single-attempt state-of-the-art (SOTA), achieved by a mixture of GPT-4o and Claude 3.5 Sonnet models, is 43% [1]. When given a single attempt, DeepSeek-Coder-V2-Instruct [20] achieves only 15.9% on the benchmark. By simply increasing the number of attempts to 250, we increase the fraction of solved problems to 56%, exceeding the single-attempt state-of-the-art by 13%.
@@
In addition to improving model quality, repeated sampling provides a new mechanism for minimizing LLM inference costs. When holding the total number of inference FLOPs constant, we find that on some datasets (e.g. MATH) coverage is maximized with a smaller model and more attempts, while on others (e.g CodeContests) it is better to use a larger model. We also compare API prices between DeepSeek-Coder-V2-Instruct, GPT-4o, and Claude Sonnet 3.5 in the context of solving SWE-bench Lite issues. When keeping the agent framework (Moatless Tools [67]) constant, sampling five times from the weaker and cheaper DeepSeek model solves more issues than single attempts from Claude or GPT, while also being over 3x cheaper.
Finally, in math word problem settings, where answers cannot be automatically verified by existing tools, we identify a large gap between coverage and the performance of common methods for deciding on a final answer. When solving MATH [26] problems with Llama-3-8B-Instruct, coverage increases from 79.8% with 100 samples to 95.3% with 10,000 samples. However, methods such as majority voting and using reward models plateau with a lower sample budget, scaling only from 38.7% to 39.8% over the same range. These results highlight that building robust verifiers remains an open problem.
@@
    We demonstrate that scaling inference compute through repeated sampling leads to large improvements in coverage across a variety tasks, models, and sample budgets. This makes it possible, and sometimes cost-effective, to amplify weaker models with many samples and outperform single attempts from more capable models. Notably, we are able to solve 56% of issues from SWE-bench Lite by sampling 250 times from DeepSeek-Coder-V2-Instruct, exceeding the single-attempt SOTA of 43%.
@@
    In domains without automatic verifiers, we show that common approaches to verification like majority voting and reward model scoring plateau beyond approximately 100 samples. This leads to a growing gap between the performance achieved with these methods and the coverage upper bound.
@@ 2.3 Repeated Sampling Can Help Balance Performance and Cost
FLOPs per token $\approx 2*\left(\text{num parameters}+2*\text{num layers}*\text{token dim}*\text{context length}\right)$
@@
[…]
@@ 3.1 Scaling Laws for Repeated Sampling
Figure 5: The relationship between coverage and the number of samples can be modelled with an exponentiated power law for most tasks and models. We highlight that some curves, such as Llama-3-8B-Instruct on MiniF2F-MATH, do not follow this trend closely.
@@ 4 Harnessing Repeated Sampling Requires Precision
So far, we have focused on measuring model coverage, characterizing the benefits of repeated sampling under the best-case scenario where we can always identify correct model samples. We now turn to the complementary problem of precision: given a collection of model samples, can we identify the correct ones? In Section 4.1, we evaluate two common verification methods (majority voting and reward model scoring) on GSM8K and MATH. Additionally, in Section 4.2, we discuss potential pitfalls when relying on unit tests to identify correct software programs.
@@ 4.1 Common Verification Methods Don’t Always Scale with the Sample Budget
Of the five tasks we evaluate, only GSM8K and MATH lack tools for automatically verifying solutions. Here, we evaluate two common approaches to deciding on a final answer: calculating a majority vote across samples and using a reward model to assign a score to each sample. We test these techniques on their ability to identify correct solutions from the 10,000 samples we generated with Llama-3-8B-Instruct and Llama-3-70B-Instruct in Section 2. We benchmark three methods:
@@
We use ArmoRM-Llama3-8B-v0.1 [57] as a reward model, which currently has the highest reasoning score on the RewardBench leaderboard [38] among open-weight models. We report our results in Figure 7 as we increase the number of samples. While success rate initially increases with the number of samples for all three methods, it plateaus around 100 samples. Meanwhile, coverage continues to increase with the number of samples and exceeds 95%.
In the case of majority voting, this success rate saturation is easily explainable. As the number of samples increases, the proportion of votes allocated to each answer stabilizes, and therefore the success rate plateaus. For some GSM8K and MATH problems, correct solutions are sampled with a probability of 1% or lower (see Figure 8), making them a minority of samples. As the number of samples increases, rare correct solutions will appear for more problems, increasing coverage but not the success rate with majority voting. In order to fully benefit from repeated sampling, sample identification methods must be able to solve these “needle-in-a-haystack” cases and identify rare, correct samples.
Given the poor performance of the reward model, it is reasonable to wonder how “hard” it is to verify a candidate solution. With GSM8K and MATH, only a sample’s final answer is used for assessing correctness, with the intermediate chains of thought being excluded. If models generated only non-sensical chains of thought before guessing a correct final answer, verification may not be any easier than solving the problem in the first place. We investigate this question by manually evaluating 105 chains of thought from correct Llama-3-8B-Instruct solutions to GSM8K problems, reporting our results in Table 2. We find that over 90% of the chains of thought that we graded are faithful, even among problems where correct answers are generated infrequently. These correct reasoning steps indicate that there is signal for a verifier to exploit when identifying correct samples. Interestingly, during this process we also identified one GSM8K problem that has an incorrect ground truth answer (see Appendix E). This incorrect GSM8K problem is also the only one that Llama-3-70B-Instruct did not generate a “correct” sample for across 10,000 attempts.

### v3

@@ Abstract
Scaling the amount of compute used to train language models has dramatically improved their capabilities. However, when it comes to inference, we often limit models to making only one attempt at a problem. Here, we explore inference compute as another axis for scaling, using the simple technique of repeatedly sampling candidate solutions from a model. Across multiple tasks and models, we observe that coverage – the fraction of problems that are solved by any generated sample – scales with the number of samples over four orders of magnitude. Interestingly, the relationship between coverage and the number of samples is often log-linear and can be modelled with an exponentiated power law, suggesting the existence of inference-time scaling laws. In domains like coding and formal proofs, where answers can be automatically verified, these increases in coverage directly translate into improved performance. When we apply repeated sampling to SWE-bench Lite, the fraction of issues solved with DeepSeek-Coder-V2-Instruct increases from 15.9% with one sample to 56% with 250 samples, outperforming the single-sample state-of-the-art of 43%. In domains without automatic verifiers, we find that common methods for picking from a sample collection (majority voting and reward models) plateau beyond several hundred samples and fail to fully scale with the sample budget.
00footnotetext: Code: https://github.com/ScalingIntelligence/large_language_monkeys.00footnotetext: Data: https://huggingface.co/datasets/ScalingIntelligence/monkey_business.00footnotetext: Equal Contribution. Work done by BB as a visiting researcher at Stanford.
@@ 1 Introduction
The ability of large language models (LLMs) to solve coding, mathematics, and other reasoning tasks has improved dramatically over the past several years [47, 11, 2, 4]. Scaling the amount of training compute through bigger models, longer pre-training runs, and larger datasets has been a consistent driver of these gains [27, 37, 28].
In contrast, a comparatively limited investment has been made in scaling the amount of computation used during inference. Larger models do require more inference compute than smaller ones, and prompting techniques like chain-of-thought [61] can increase answer quality at the cost of longer (and therefore more computationally expensive) outputs. However, when interacting with LLMs, users and developers often restrict models to making only one attempt when solving a problem.
In this work, we explore repeated sampling (Figure 1) as a simple approach to scaling inference compute in order to improve reasoning performance. Existing work provides encouraging examples that repeated sampling can be beneficial in math, coding, and puzzle-solving settings [60, 48, 23]. Notably, AlphaCode [41], a state-of-the-art system for competitive programming, finds that performance continues to improve with a million samples per problem. Our goal is to systematically characterize these benefits across a range of tasks, models, and sample budgets.
Figure 1: The repeated sampling procedure that we follow in this paper. 1) We generate many independent candidate solutions for a given problem by sampling from an LLM with a positive temperature. 2) We use a domain-specific verifier (ex. unit tests for code) to select a final answer from the generated samples.
@@
    Precision: How often can we identify correct samples from our collection of generations?
Both properties are needed for achieving strong real-world performance. With unlimited samples, any model that assigns a non-zero probability to every sequence will achieve perfect coverage. However, repeated sampling is only practical if we can improve coverage with a feasible budget. Similarly, generating large sample collections is only useful if the correct samples in a collection can be identified. The difficulty of the precision problem can vary by task. In some settings, existing tools like proof checkers and unit tests can automatically verify every sample. In other cases, like when solving word problems, other methods for verification are needed.
@@
Exploring coverage first, we find that sampling up to 10,000 times per problem can significantly boost coverage on math and coding tasks (Section 2). When solving CodeContests [41] programming problems using Gemma-2B [52], we increase coverage by over 300x, from 0.02% with one sample to 7.1% with 10,000 samples. Interestingly, the relationship between $\log(\text{coverage})$ and the number of samples often follows an approximate power law (Section  3). With Llama-3 [3] and Gemma models, this leads to coverage growing nearly log-linearly with the number of samples over several orders of magnitude.
@@
In settings with automatic verification tools, increases in coverage translate directly into improved task performance. When applying repeated sampling to competitive programming and writing Lean proofs, models like Llama-3-8B-Instruct can exceed the single-sample performance of much stronger ones like GPT-4o [2]. This ability to amplify weaker models extends to the challenging SWE-bench Lite dataset of real-life GitHub issues [32], where the current single-sample state-of-the-art (SOTA), achieved by a mixture of GPT-4o and Claude 3.5 Sonnet, is 43% [1]. When restricted to a single sample, DeepSeek-Coder-V2-Instruct [20] solves only 15.9% of issues. By simply increasing the number of samples to 250, we increase the fraction of solved issues to 56%, exceeding the state-of-the-art by 13%.
@@
In addition to improving model quality, repeated sampling provides a new mechanism for minimizing LLM inference costs (Section 2.3). When holding the total number of inference FLOPs constant, we find that on some datasets (e.g. MATH), coverage is maximized with a smaller model and more samples, while on others (e.g CodeContests) it is better to sample fewer times from a larger model. We also compare API prices between DeepSeek-Coder-V2-Instruct, GPT-4o, and Claude Sonnet 3.5 in the context of solving SWE-bench Lite issues. When keeping the agent framework (Moatless Tools [67]) constant, sampling five times from the weaker and cheaper DeepSeek model solves more issues than single samples from Claude or GPT while also being over 3x cheaper.
Finally, we demonstrate that scalable verification is necessary for fully benefiting from repeated sampling. As the number of samples increases, coverage improves through models generating correct solutions to problems they have not previously solved. However, these increasingly rare correct generations are only beneficial if verifiers can “find the needle in the haystack” and identify them from collections of mostly-incorrect samples. In math word problem settings, we find that two common methods for verification (majority voting and reward models) do not possess this ability. When solving MATH [26] problems with Llama-3-8B-Instruct, coverage increases from 82.9% with 100 samples to 98.44% with 10,000 samples. However, when using majority voting or reward models to select final answers, the biggest performance increase is only from 40.50% to 41.41% over the same sample range. As the number of samples increases, the gap between coverage (i.e. performance with a perfect verifier) and the performance of these methods increases as well (Figure  7).
@@
    We demonstrate that scaling inference compute through repeated sampling leads to large improvements in coverage across a variety of tasks and models. This makes it possible, and sometimes cost-effective, to amplify weaker models with many samples and outperform single samples from more capable models.
@@
    In domains without automatic verifiers, we show that common approaches to verification plateau beyond approximately 100 samples. This leads to a growing gap between the performance achieved with these methods and the coverage upper bound.
@@ 2.3 Repeated Sampling Can Help Balance Performance and Cost
TotalInferenceFLOPs $\approx\left(\sum_{t=1}^{\text{NumPromptTokens}}\text{FLOPsPerToken}(t)\right)+$
$\left(\sum_{t=1}^{\text{NumDecodeTokens}}\text{FLOPsPerToken}(t+\text{NumPromptTokens})*\text{NumCompletions}\right)$
@@ 3.1 Scaling Laws for Repeated Sampling
Figure 5: The relationship between coverage and the number of samples can be modelled with an exponentiated power law for most tasks and models. We highlight that some curves, such as Llama-3-8B-Instruct on MiniF2F-MATH, do not follow this trend closely. We show the mean and standard deviation of the error between the coverage curve and the power law fit across 100 evenly sampled points on the log scale.
@@ 4 Harnessing Repeated Sampling Requires Precision
So far, we have focused on measuring model coverage, characterizing the benefits of repeated sampling under the scenario where we can always identify correct model samples. We now turn to the complementary problem of precision: given a collection of model samples, how often can we identify the correct ones? In particular, we are interested in the performance of verifiers as we scale up the number of samples. For some problems, correct solutions are sampled from the model at low probabilities (e.g. 1% or lower, see Figure 8). As the number of samples increases and rare, correct solutions are generated for more problems, model coverage improves. In order to convert these coverage improvements into higher success rates, verifiers must be able to find the “needle in the haystack” and identify infrequent correct samples.
@@ 4.1 Common Verification Methods Don’t Always Scale with the Sample Budget
Of the five tasks we evaluate, only GSM8K and MATH lack tools for automatically verifying solutions. We test three simple and commonly used verification approaches on their ability to identify correct solutions from these datasets:
@@
We reuse the collections of 10,000 samples that we generated with Llama-3-8B-Instruct and Llama-3-70B-Instruct in Section 2. We use ArmoRM-Llama3-8B-v0.1 [57] as a reward model, which scores highly on the reasoning section of the RewardBench leaderboard [39]. We report our results in Figure 7 as we increase the number of samples. While success rates initially increase with the number of samples for all three methods, they plateau around 100 samples. Meanwhile, coverage continues to increase with the number of samples and eventually exceeds 95%. In the case of majority voting, this success rate saturation is intuitive, since the occurrence of rare, correct solutions does not affect the most common answer that majority voting chooses.
Given the poor performance of these verifiers (in particular the reward model), it is reasonable to wonder how “hard” it is to verify a candidate solution. With GSM8K and MATH, only a sample’s final answer is used for assessing correctness, with the intermediate chains of thought being discarded. If models generated only non-sensical chains of thought before guessing a correct final answer, verification may not be any easier than solving the problem in the first place. We investigate this question by manually evaluating 105 chains-of-thought from correct Llama-3-8B-Instruct samples to GSM8K problems, reporting our results in Table 2.
We find that over 90% of the chains-of-thought that we graded are faithful, even among problems where correct answers are generated infrequently. These correct reasoning steps indicate that there is signal for a verifier to exploit when identifying correct samples. Interestingly, during this process we also identified one GSM8K problem that has an incorrect ground truth answer (see Appendix E). This incorrect GSM8K problem is also the only one that Llama-3-70B-Instruct did not generate a “correct” sample for across 10,000 attempts.